OpenStudy (anonymous):

a boy uns 500m in 1min40s, which is the speed in m/s?

OpenStudy (anonymous):

speed =distance/time =500/(1x60+40) =5 m/s

OpenStudy (anonymous):

speed = distance travelled / time = 500/ (60+40) <--- convert to 1 min to 60s so total time =60+40 =5 m/s